# An Effective Incorporating Heterogeneous Knowledge Curriculum Learning for Sequence Labeling

### v1

@@ Abstract
Sequence labeling models often benefit from incorporating external knowledge. However, this practice introduces data heterogeneity and complicates the model with additional modules, leading to increased expenses for training a high-performing model. To address this challenge, we propose a two-stage curriculum learning (TCL) framework specifically designed for sequence labeling tasks. The TCL framework enhances training by gradually introducing data instances from easy to hard, aiming to improve both performance and training speed. Furthermore, we explore different metrics for assessing the difficulty levels of sequence labeling tasks. Through extensive experimentation on six Chinese word segmentation (CWS) and Part-of-speech tagging (POS) datasets, we demonstrate the effectiveness of our model in enhancing the performance of sequence labeling models. Additionally, our analysis indicates that TCL accelerates training and alleviates the slow training problem associated with complex models.
Keywords: Curriculum Learning, Sequence Labeling, Difficulty Measure
@@
[…]
${}^{1}$ Department of Information Management, Peking University
${}^{2}$ Research Center for Digital Humanities of Peking University
${}^{3}$ School of Foreign Languages, Peking University
tangxuemei@stu.pku.edu.cn
Abstract content
1.   Introduction
Sequence labeling is a fundamental task in natural language processing (NLP) that involves assigning labels to individual elements in a sequence. Recent advancements in neural network approaches have demonstrated remarkable performance in sequence labeling tasks  Zhang et al. (2014); Chen et al. (2017); Zhang et al. (2018); Tian et al. (2020a); Nguyen et al. (2021); Hou et al. (2021); Liu et al. (2021). Some of these works have explored the integration of external knowledge, such as n-grams, lexicons, and syntax, to enhance sequence labeling models. However, this integration increases the heterogeneity and complexity of the input data. Furthermore, incorporating knowledge into sequence labeling models often requires additional encode modules, such as attention mechanisms  Liu et al. (2021); Tian et al. (2020a) or graph neural network (GNN) Gui et al. (2019); Nie et al. (2022). These additional modules increase the model’s parameters, making it more time-consuming and resource-intensive to develop a high-performing system.
Curriculum Learning (CL) is an effective approach to tackle these challenges  Bengio et al. (2009), as it mimics the human learning process by gradually introducing training samples from easy to hard, thereby enabling efficient learning in heterogeneous data while improving both the speed and quality of the model Wang et al. (2021). CL has demonstrated success in various NLP tasks, including machine translation Wan et al. (2020), dialogue generation Zhu et al. (2021), and text classification Zhang et al. (2022).
In CL, data-selection curriculum strategies play a crucial role. Recently,  Mohiuddin et al. (2022) proposed metrics that rely on the similarity of parallel sentence pairs to guide data selection in machine translation tasks. Similarly,  Yuan et al. (2022) introduced difficulty metrics for entity typing based on training loss. However, these difficulty metrics primarily focus on the sentence level, and there is a lack of token-level and word-level metrics to measure the difficulty of sequence labeling tasks.
To address this gap, in this paper, we introduce a two-stage curriculum learning (TCL) framework specifically designed for sequence labeling tasks. The first stage is data-level CL, where we train a basic teacher model on all available training data, aiming to alleviate the cold start problem of the student model. The second stage is model-level CL, where we start training the student model on a selected subset of the teacher model and gradually expand the training subset by considering the difficulty of the data and the state of the student model. Furthermore, we explore different difficulty metrics for sequence labeling tasks within the TCL framework. These metrics include a pre-defined metric, such as sentence length, and model-aware metrics, namely Top-N least confidence (TLC), Maximum normalized log-probability (MNLP), and Bayesian uncertainty (BU). Finally, we choose the classical sequence labeling tasks, Chinese word segmentation and part-of-speech tagging, to validate our proposed approach.
In summary, our work makes the following contributions:
  • We propose a TCL framework for sequence labeling tasks. In the first stage, we train a simple transfer teacher model using all available data, which establishes the initial sample order for the student model and helps it warm up. In the second stage, we dynamically adjust the order of samples based on the current state of the student model.
  • We design and compare various difficulty metrics tailored for sequence labeling tasks within the TCL framework. These metrics offer a more comprehensive assessment of the difficulty at the token and word levels, going beyond the traditional sentence-level metrics.
  • We conduct extensive evaluations on six datasets encompassing sequence labeling tasks. Experiments and analyses show that our model accelerates training speed and the proposed difficulty metrics can plan more efficient learning paths for the model. Moreover, our approach exhibits flexibility and can be easily applied to other sequence labeling models.
2.   Related Work
In recent years, neural network methods have achieved remarkable success in sequence labeling tasks Zhang et al. (2014); Chen et al. (2017); Zhang et al. (2018); Gui et al. (2019); Tian et al. (2020a); Nguyen et al. (2021); Hou et al. (2021); Lobov et al. (2022). These studies commonly employ the fusion of external knowledge to enhance the performance of sequence labeling models. This external knowledge includes pre-trained embeddings Zhang et al. (2018); Shao et al. (2017), n-grams and lexicons Tian et al. (2020b); Liu et al. (2021), and syntax information Tian et al. (2020a); Tang et al. (2022). Specialized structures are employed to incorporate the knowledge, such as various variants of attention modules Liu et al. (2021); Tian et al. (2020a); Tang and Su (2022) and graph neural networks Gui et al. (2019); Huang et al. (2021); Tang et al. (2022). These structures effectively encode the relevant information from the input data. Nevertheless, their integration poses two challenges. Firstly, the inclusion of words, n-grams, and syntax information introduces heterogeneity and complexity to the input data. Secondly, the encoding modules increase the number of training parameters, making it more computationally expensive to train a model.
Curriculum learning (CL), introduced by Bengio et al. (2009), revolves around the principle of “training from easier data to harder data”. CL has been successfully applied to various NLP tasks, including machine translation Wan et al. (2020), dialogue generation Zhu et al. (2021), and text classification Zhang et al. (2022). It serves as an effective approach to address data heterogeneity and improve the performance of tasks.
The CL framework comprises two key components: difficulty measurement and training schedule. The former determines the relative “easiness” of each data example, which can be pre-defined or determined automatically. The latter determines the sequence of data subsets during the training process,guided by the judgment from the difficulty measurer Wang et al. (2021), using either a linear or non-linear scheduler. When both components are designed based on human prior knowledge, it is referred to as pre-defined CL. On the other hand, if one of the components is driven by a data-driven algorithm, it is known as automatic CL, exemplified by self-paced learning and transfer teacher. Pre-defined difficulty measurer is typically designed by considering data characteristics such as sentence length and word rarity. In general, designing difficulty metrics specifically tailored to the task at hand is more efficient and typically involves data-driven algorithms. For instance,  Mohiuddin et al. (2022) developed three deterministic scoring methods for machine translation that leverage the similarity of a parallel sentence pairs and cross-entropy.  Zhu et al. (2021) evaluated sample difficulty in dialogue generation based on model loss.  Lobov et al. (2022) utilized metrics such as sentence length, average confidence, and perplexity value to assess difficulty in named entity recognition (NER). However, these metrics primarily concentrate on the sentence-level and do not adequately address token- and word-level difficulty.
3.   Method
The framework proposed in this study consists of three main components: a teacher sequence labeling model, a student sequence labeling model, and a TCL training strategy. It is worth noting that the TCL is independent of the sequence labeling model. In line with previous studies, we adopt an encoder-to-decoder framework for the sequence labeling model, where the encoder is typically a pre-trained language model (PLM), and the decoder can be either a conditional random field (CRF) layer or a Softmax layer. Figure 1 illustrates the overall architecture of the proposed model.
Figure 1: The framework of the proposed model consists of a teacher model, a student model, and a TCL strategy. Here, “Root f” represents Root function.
3.1.   Two-stage Curriculum Learning
Algorithm 1 Training Process
1:Original corpus D, difficulty metric $S(\cdot)$, training epochs of the teacher model $E_{0}$, training epochs of the student model $E_{s}$, training scheduler $\lambda$, length function ||$\cdot$|| ⋅ |.
2:A trained student model $\theta$
3:// data level curriculum learning
4:Train the teacher model $\theta_{0}$ with D for $E_{0}$ epochs
5:Get $S(\theta_{0})$ for each sample in D
[…]
7:// model level curriculum learning
8:Initial size of the training set of the student model: $m=\lambda|D|$
9:Initial training set for the student model: $D_{s}=D_{r}[0:m]$
10:The remaining training data: ${D}_{o}={D}_{r}[m:]$
11:for $epoch=1,2,..E_{s}$ do
12:     if  $\lambda$ <1 then
[…]
14:         current model $\theta_{*}$
[…]
18:         d) Update $\lambda$
19:         e) Obtain the number of new training
20:         data: $m=\lambda|D|-|D_{s}|$
21:         f) Add new training data into the
22:         student training set: ${D}_{s}+={D}_{r}[0:m]$
23:         g) The remaining data: ${D}_{o}={D}_{r}[m:]$
24:     else
[…]
26:     end if
27:end for
We present a novel two-stage curriculum learning approach: data-level CL and model-level CL. The Algorithm 1 provides a detailed explanation of the process.
At the data level, the algorithm begins by training a basic transfer teacher model on the entire training dataset $D$ for a specified number of epochs ($E_{0}$) to obtain the teacher model parameter $\theta_{0}$ (Line 1). It is worth noting that $E_{0}$ is selected to be smaller than the total number of epochs required for achieving convergence.
[…]
At the model level, the algorithm addresses the cold-start problem by initializing the training set of the student model, denoted as $D_{s}$, with a certain percentage of samples from $D_{r}$ (Line 4-5). The parameter $\lambda$ represents the proportion of these samples and plays a crucial role in controlling the curriculum learning process. The remaining data in $D_{r}$ that is not included in $D_{s}$ is represented as $D_{o}$ (Line 6). Gradually, the algorithm incorporates the samples from $D_{o}$ into $D_{s}$ by adjusting the threshold parameter $\lambda$. In the algorithm, we denote the number of new data to be added by $m$, and the value of $m$ is determined by $\lambda$ (Line 4, 15-16).
Next, we describe the update process of $D_{s}$. In the algorithm, the student model is initially trained on $D_{s}$ to obtain the updated model parameter $\theta_{*}$ (Line 9-10). The remaining data $D_{o}$ is then ranked according to the difficulty metric $S(\theta)$, resulting in a new ranked dataset $D_{r}$ (Line 11-13).
The algorithm proceeds to update the threshold parameter $\lambda$ (line 14) and calculates the number of new training samples to be added to the student training set $D_{s}$ from $D_{r}$ based on this parameter (lines 15-19). This process gradually incorporates new and more challenging samples into the student model training set. Consequently, as $\lambda$ reaches 1, all the data in $D_{o}$ is incorporated into $D_{s}$. Finally, the complete dataset is used to train the student model until convergence is achieved.
In algorithm 1, two crucial components play a significant role: the difficulty metrics $S(\cdot)$ and the threshold $\lambda$. These elements control the difficulty ranking of training samples and the progression of training samples feeding into the model, respectively. In the following sections, we will describe the design of each of these two elements.
3.2.   Difficulty Metrics
In sequence labeling tasks, the difficulty of samples is intricately associated with each token in the sequence. However, since it is difficult to assess the difficulty of individual tokens, we draw on the concept of uncertainty commonly used in active learning to reflect the difficulty of the samples, which quantifies whether the current model is confident or hesitant in labeling the training samples. In short, if the model is confident in predicting a sample, it is considered an easy sample. Therefore, the difficulty of the sample is dynamically evaluated based on the model capacity.
Next, we provide detailed formulations of various strategies for computing the difficulty $S(\cdot)$ in Algorithm 1.

### v2

@@ Abstract
Sequence labeling models often benefit from incorporating external knowledge. However, this practice introduces data heterogeneity and complicates the model with additional modules, leading to increased expenses for training a high-performing model. To address this challenge, we propose a dual-stage curriculum learning (DCL) framework specifically designed for sequence labeling tasks. The DCL framework enhances training by gradually introducing data instances from easy to hard. Additionally, we introduce a dynamic metric for evaluating the difficulty levels of sequence labeling tasks. Experiments on several sequence labeling datasets show that our model enhances performance and accelerates training, mitigating the slow training issue of complex models  [^1] https://github.com/tangxuemei1995/DCL4SeqLabeling.
Xuemei Tang1, Jun Wang2, Qi Su3thanks: Corresponding authors., Chu-Ren Huang1, and Jinghang Gu [^1] footnotemark: 1 1The Hong Kong Polytechnic University, Hong Kong, China 2Department of Information Management, Peking University, Beijing, China 3School of Foreign Languages, Peking University, Beijing, China xuemeitang00@gmail.com,  {sukia,junwang}@pku.edu.cn {churen.huang, jinghang.gu}@polyu.edu.hk
1 Introduction and Related Work
Sequence labeling is a core task in natural language processing (NLP) that involves assigning labels to individual elements in a sequence. Recent advancements in neural network methods have significantly improved performance in sequence labeling tasks  Zhang et al. (2014); Chen et al. (2017a); Zhang et al. (2018); Tian et al. (2020a); Nguyen et al. (2021); Hou et al. (2021); Liu et al. (2021). Some studies have explored integrating external knowledge, such as n-grams, lexicons, and syntax, to enhance these models. However, this integration adds heterogeneity and complexity to the input data. Additionally, incorporating such knowledge often necessitates extra encoding modules, like attention mechanisms  Liu et al. (2021); Tian et al. (2020b) or graph neural networks (GNN) Chen et al. (2017b); Gui et al. (2019); Nie et al. (2022), which increase model parameters and make the system more computationally expensive to develop.
Curriculum Learning (CL) (Bengio et al., 2009) effectively addresses these challenges by simulating the human learning process, where training samples are introduced progressively from easy to hard. This approach facilitates efficient learning from heterogeneous data while enhancing both the speed and performance of the model Bengio et al. (2009); Wang et al. (2021). CL has shown success in a variety of NLP tasks, including machine translation Wan et al. (2020), dialogue generation Zhu et al. (2021), and text classification Zhang et al. (2022). Data-selection strategies are crucial in CL. However, these difficulty metrics primarily focus on the sentence level, such as Mohiuddin et al. (2022),  Yuan et al. (2022) and  Liu et al. (2024)’s works, and there is a lack of token-level and word-level metrics to measure the difficulty of sequence labeling tasks.
To address this gap, in this paper, we introduce a dual-stage curriculum learning (DCL) framework specifically designed for sequence labeling tasks. The first stage is data-level CL, where we train a basic teacher model on all available training data, aiming to alleviate the cold start problem of the student model. The second stage is model-level CL, where we start training the student model on a selected subset of the teacher model and gradually expand the training subset by considering the difficulty of the data and the state of the student model. Furthermore, we explore different difficulty metrics for sequence labeling tasks within the DCL framework. These metrics include a pre-defined metric, such as sentence length, and model-aware metrics, namely Top-N least confidence (TLC), Maximum normalized log-probability (MNLP), and Bayesian uncertainty (BU). Finally, we choose the classical sequence labeling tasks, Chinese word segmentation (CWS), part-of-speech (POS) tagging, and named entity recognition (NER), to validate our proposed approach.
2 Method
The framework proposed in this study consists of three main components: a teacher sequence labeling model, a student sequence labeling model, and a DCL training strategy. It is worth noting that the DCL is independent of the sequence labeling model.
Figure 1: The framework of the proposed model consists of a teacher model, a student model, and a DCL strategy. Here, “Root f” represents Root function.
2.1 Dual-stage Curriculum Learning
Algorithm 1 Training Process with DCL
1:Original corpus $\mathcal{D}$, difficulty metric $S(\cdot)$, teacher model epochs $E_{0}$, student model epochs $E_{s}$, scheduler $\lambda$, length function ||$\cdot$|| ⋅ |
2:Trained student model $\theta$
3:// Data-level Curriculum Learning
4:Train teacher model $\theta_{0}$ on $\mathcal{D}$ for $E_{0}$ epochs
5:Compute $S(\theta_{0})$ for each sample in $\mathcal{D}$
6:Sort $\mathcal{D}$ by $S(\theta_{0})$ in ascending order to obtain ranked dataset $\mathcal{D}_{r}$
7:// Model-level Curriculum Learning
8:Initialize $\lambda_{0}$ (starting curriculum ratio)
9:$m\leftarrow\lambda_{0}\cdot|\mathcal{D}|$
10:Student training set $\mathcal{D}_{s}\leftarrow\mathcal{D}_{r}[0:m]$
11:Remaining data $\mathcal{D}_{o}\leftarrow\mathcal{D}_{r}[m:]$
12:for $\textit{epoch}=1$ to $E_{s}$ do
13:     if $\lambda<1$ then
14:         a) Train student model on $\mathcal{D}_{s}$ to obtain current $\theta_{*}$
15:         b) Compute $S(\theta_{*})$ for all samples in $\mathcal{D}_{o}$
16:         c) Sort $\mathcal{D}_{o}$ by $S(\theta_{*})$ in ascending order to get updated $\mathcal{D}_{r}$
17:         d) Update $\lambda$ using Eq. 6
18:         e) Calculate new data size: $m\leftarrow\lambda\cdot|\mathcal{D}|-|\mathcal{D}_{s}|$
19:         f) Expand $\mathcal{D}_{s}$ with new samples: $\mathcal{D}_{s}\mathrel{+}=\mathcal{D}_{r}[0:m]$
20:         g) Update remaining data: $\mathcal{D}_{o}\leftarrow\mathcal{D}_{r}[m:]$
21:     else
22:         Train student model on $\mathcal{D}_{s}$
23:     end if
24:end for
We propose a novel dual-stage curriculum learning approach: data-level CL and model-level CL, as detailed in Algorithm 1.
At the data level, we first train a basic teacher model on the entire dataset $\mathcal{D}$ for $E_{0}$ epochs, where $E_{0}$ is smaller than the total epochs needed for convergence (Line 1). The teacher model $\theta_{0}$ is then used to calculate difficulty scores $S(\theta_{0})$ for each sample (Line 2), and the samples are sorted by difficulty to form a ranked dataset $\mathcal{D}_{r}$ (Line 3).
At the model level, we address the cold-start issue by initializing the student model training set $\mathcal{D}_{s}$ with a subset of $\mathcal{D}_{r}$ (Lines 4-6). The proportion of samples, controlled by the parameter $\lambda$, governs the curriculum learning process. The remaining data, $\mathcal{D}_{o}$, is incorporated into $\mathcal{D}_{s}$ gradually as $\lambda$ increases. The number of new samples to be added is denoted as $m$ (Lines 5, 14).
The student model is trained on $\mathcal{D}_{s}$ to update the model parameters $\theta_{*}$ (Line 10). Then, $\theta_{*}$ is used for the difficulty calculation of the samples in $\mathcal{D}_{o}$ (Line 11). Next, $\mathcal{D}_{o}$ is ranked by new difficulty scores, forming a new ranked dataset $\mathcal{D}_{r}$ (Line 12). The threshold $\lambda$ is updated (Line 13), and new samples are added to $\mathcal{D}_{s}$ based on $\lambda$ (Lines 14-15). As $\lambda$ approaches 1, all of $\mathcal{D}_{o}$ is added to $\mathcal{D}_{s}$. The complete dataset is then used to train the student model to convergence.
The key elements in Algorithm 1 are the difficulty metric $S(\cdot)$ and threshold $\lambda$, which control the difficulty ranking of samples and the progression of training, respectively. The design of these components will be discussed in the following sections.
2.2 Difficulty Metrics
We now provide a detailed formulation for calculating the difficulty $S(\cdot)$ in Algorithm 1. In sequence labeling tasks, sample difficulty is tied to individual tokens, but assessing token-level difficulty is challenging. We use uncertainty from active learning to measure the model’s confidence in labeling training samples.
Bayesian Uncertainty (BU). Following Buntine and Weigend (1991), model uncertainty can be assessed using Bayesian Neural Networks. As noted by Wang et al. (2019), higher predicted probability variance indicates greater uncertainty, suggesting that the model is less confident about the sample. In this work, we employ the widely-used Monte Carlo dropout Gal and Ghahramani (2016) to approximate Bayesian inference.
First, we apply Monte Carlo dropout Gal and Ghahramani (2016) to obtain each sample of token-level tagging probabilities. Specifically, for each token $x_{i}$, we perform $K$ stochastic forward passes through the model, resulting in $K$ predicted distributions ${P(y_{i}\mid x_{i})}_{1},\ldots,{P(y_{i}\mid x_{i})}_{K}$. This provides $K$ predictions with associated probabilities for each token. Then the expectation of token-level tagging probability can be approximated by
$\mathbb{E}[P(y_{i}|x_{i})]\approx\frac{1}{K}\sum_{k=1}^{K}{P(y_{i}|x_{i})}_{k}$
$var(x_{i},\theta)\approx$ (2)
@@ 2.2 Difficulty Metrics
$var(\theta)_{aver.}=\frac{1}{M}\sum_{i=1}^{M}var(x_{i},\theta)$ (3)
$var(\theta)_{max}=max_{i\in[1,M]}var(x_{i},\theta)$ (4)
$S(\theta)^{BU}=var(\theta)_{max}+var(\theta)_{aver.}$ (5)
2.3 Training Scheduler
The training scheduler regulates the pace of CL. In our approach, we employ the Root function as the control mechanism. This function ensures that the model receives sufficient time to learn newly introduced examples while gradually decreasing the number of newly added examples throughout the training process.
$\lambda=\min\left(1,\sqrt{\frac{1-\lambda_{0}^{2}}{E_{grow}}\cdot t+\lambda_{0}^{2}}\right)$ (6)
where $E_{grow}$ denotes the number of epochs required for $\lambda$ to reach 1, while $\lambda_{0}>0$ represents the initial proportion of the easiest training samples. $t$ indicates the $t_{th}$ training epochs. When $\lambda$ reaches 1, the model has access to the entire training dataset.
 Model CL Setting CTB5 CTB6 PKU
CWS POS CWS POS CWS POS
  Tian et al. (2020a) - 98.73 96.60 97.30 94.74 - -
Tian et al. (2020b) (McASP) - 98.77 96.77 97.43 94.82 - -
Liu et al. (2021) - - 97.14 - - - -
Tang et al. (2024) (SynSemGCN) - 98.83 96.77 97.86 94.98 98.05 95.50
McASP Rand. 98.81 96.84 97.37 94.90 98.38 96.27
Length 98.83 96.85 97.35 94.82 98.40 96.25
TLC 98.83 96.89 97.37 94.83 98.41 96.30
MNLP 98.85 96.81 97.41 94.92 98.41 96.30
BU 98.91 96.87 97.42 94.90 98.43 96.32
SynSemGCN Rand. 98.84 97.86 97.99 95.05 98.48 96.40
Length 98.80 96.84 97.40 94.94 98.53 96.48
TLC 98.83 97.81 97.98 95.02 98.61 96.55
MNLP 98.78 97.72 98.04 95.13 98.56 96.48
BU 98.90 97.95 98.05 95.14 98.59 96.54
Table 1: Experimental results of different models using different CL settings on test sets of three datasets. Here, “CWS” represents the F1 value of CWS, and “POS” means the F1 value of the joint CWS and POS tagging. “-” means without the CL training strategy, and “TLC”, “MNLP”, and “BU” means using the DCL setting with different difficulty metrics. The maximum F1 scores for each dataset are highlighted.
3 Experiments
3.1 Dataset and Experimental Configurations
Dataset. Chinese word segmentation (CWS) and part-of-speech (POS) tagging are representative sequence labeling tasks. So we evaluate our approach using three CWS and POS tagging datasets, including Chinese Penn Treebank version 5.0  [^2] https://catalog.ldc.upenn.edu/LDC2005T01, 6.0  [^3] https://catalog.ldc.upenn.edu/LDC2007T36, and PKU. More dataset details can be found in Appendix A.
Teacher and student models. In this study, the basic transfer teacher framework is RoBERTa + Softmax. For the student model, we select two representative complex models introduced by Tian et al. (2020b) and Tang et al. (2024). In their work, Tian et al. (2020b) employed an attention mechanism framework, McASP, to integrate lexicons and n-grams for the joint CWS and POS tagging task, using BERT as the encoder. Meanwhile, Tang et al. (2024) incorporated syntax and semantic knowledge into sequence labeling tasks through a GCN framework called SynSemGCN, with RoBERTa as the sequence encoder.
Curriculum learning baselines. We compare our difficulty metric with four baseline difficulty metrics for CL: a. Random: Samples are assigned in random order; b. Sentence Length (Length): Samples are ranked from shortest to longest, based on the intuition that longer sequences are more challenging to encode; (Random and Length metics represent simple CL, namely without the teacher model). c. Top-N Least Confidence (TLC): The difficulty of a sequence is determined by using the $N$ tokens with the lowest confidence; d. Maximum Normalized Log-Probability (MNLP): The difficulty is assessed by calculating the product of the label probabilities for all tokens in the sequence. The detailed computation processes for TLC and MNLP are provided in Appendix B.
For further details on the important hyper-parameters of the model, please refer to Appendix C. We discuss the selection process of these parameter values in detail in the Appendix D.
3.2 Overall Experimental Results
Table 1 presents the experimental results of baselines and two models with different CL settings. The experimental results reveal several noteworthy conclusions.
Firstly, the DCL methodology introduced in this paper is flexible and can be integrated with various complex models. As shown in Table 1, the difficulty metrics proposed here outperform the Random and Length metrics across most datasets. Specifically, the BU metric consistently delivers the best performance on the majority of datasets when applied to the SynSemGCN model, surpassing the TLC and MNLP metrics.
Additionally, we compare our approach with previous methods that incorporate external knowledge or resources into the encoder. The results reveal that models using CL exhibit significant performance improvements, surpassing the performance of earlier methods.
     Model CTB5 Time
CWS POS
Ours 98.90 97.95 287m
w/o data CL(BU) 98.90 97.88 -
w/o model CL(BU) 98.85 97.51 -
w/o DCL 98.75 96.73 393m
Table 2: Ablation experimental results of DCL. The baseline model “w/o DCL” denotes the model SynSemGCN; “w/o model CL” means the student model always uses the initial data order sorted by the transfer teacher model; “w/o data CL” indicates the initial training samples for the student model is drawn randomly from the training set; “Ours” indicates “SynSemGCN+DCL(BU)”. Both teacher and student models with DCL in this table use BU as the difficulty metric. “Time” means the training time (in minutes).
  Model CTB5
CWS POS
McASP with BU 98.91 96.87
 w/o $var(\theta)_{max}$ 98.78 96.78
 w/o $var(\theta)_{aver.}$ 98.86 96.74
McASP 98.73 96.60
Table 3: Ablation experimental results of two parts in BU metrics (Eq. 5).
Models Weibo (Chinese) Note4 (Chinese) CoNLL-2003 (English)
BERT 66.22 79.15 90.94
BERT + CL (Length) 66.81 79.63 90.79
BERT + DCL (TLC) 67.52 79.53 91.30
BERT + DCL (MNLP) 65.73 79.95 91.15
BERT + DCL (BU) 66.74 80.02 91.77
Table 4: Performance comparison of different difficulty metrics on three NER datasets.
3.3 Effect of Dual-stage Curriculum Learning
In this section, we discuss the impact of DCL. We perform ablation studies by removing either the data-level CL or the model-level CL. The results are summarized in Table 2. Model-level CL has a more significant impact than data-level CL. This is intuitive, as model-level CL influences the entire training process, while data-level CL primarily affects the early stages of student model training.
We also compare the training time of models with and without DCL. The experimental results in Table 2 show that all models were trained for 50 epochs. The training time for models using DCL includes the time spent on training the teacher model and calculating the difficulty values for the student model. The results indicate that DCL improves model performance and reduces training time by over 25%.
3.4 Ablation Study on BU Difficulty Components
We adopt McASP Tian et al. (2020b) as the backbone model, incorporating DCL as the training strategy and BU as the difficulty metric. To examine the contribution of each component in BU, we conduct ablation experiments on its two parts: $var(\theta)_{max}$ and $var(\theta)_{aver.}$, as shown in Table 3. Removing either component results in performance degradation, indicating that both components are crucial. Moreover, the comparable drop in performance suggests that $var(\theta)_{max}$ and $var(\theta)_{aver.}$ contribute similarly to the effectiveness of DCL.
3.5 Comparison of Difficulty Metrics
In this section, we examine the impact of different difficulty metrics during the model-level CL training process for the SynSemGCN model. Figure 2 shows the F1 score change on the PKU dataset development set over the first 10 epochs of model-level CL training. After 10 epochs, all training data are used, so the initial 10 epochs highlight the effect of different metrics. From the figure, we observe that BU, in particular, achieves the best performance, indicating that uncertainty-based metrics can select samples that better align with the model’s learning trajectory, leading to faster learning.
3.6 Generalization Capability
We conduct additional experiments to demonstrate the applicability of our method to the NER task. We select two Chinese and one English NER datasets: Weibo  [^4] https://catalog.ldc.upenn.edu/LDC2013T19/, OntoNotes4  [^5] https://github.com/cchen-nlp/weiboNER and CoNLL-2003 Tjong Kim Sang and De Meulder (2003). The statistics of three datasets are shown in Table 6. We compare the performance of models using DCL and CL (Length) with a model without CL on these datasets. As shown in Table 4, the results of the DCL method outperform those of BERT+CL (Length) and BERT (no CL), indicating the effectiveness of our method. This also suggests that our method can be applied to sequence labeling tasks beyond CWS and POS tagging.
4 Conclusion
This paper introduces a novel dual-stage curriculum learning framework aimed at enhancing performance and accelerating the training process for sequence labeling tasks. Focusing on the sequence labeling task of CWS, POS tagging, and NER, this framework demonstrates its effectiveness.
Limitations
There are several limitations to our study. First, the design of our difficulty metrics involves the tuning of multiple hyperparameters, which may complicate optimization. Second, we did not explore a curriculum learning process that progresses from hard to easy examples. Third, we focused on a single variation of the $\lambda$ parameter to control CL and did not investigate alternative methods for adding training data.
Acknowledgments
This research is supported by the NSFC project “The Construction of the Knowledge Graph for the History of Chinese Confucianism” (Grant No. 72010107003) and The Hong Kong Polytechnic University Project “Evaluating the Syntax-Semantics Knowledge in Large Language Models” (Grant No. P0055270).
Appendix A Dataset
The details of the three datasets are given in Table 5. Regarding the CTB datasets, we follow the same approach as previous works Shao et al. (2017); Tian et al. (2020a) by splitting the data into train/dev/test sets. In the case of PKU, We randomly select 10% of the training data to create the development set.
                          Datasets CTB5 CTB6 PKU
Train Sent. 18k 23K 17k
Word 494k 99k 482k
Dev Sent. 350 2K 1.9k
Word 7k 60K 53k
Test Sent. 348 3K 3.6k
Word 8k 12k 97k
Table 5: Detail of the three datasets.
  Datasets Type Train Dev Test
Weibo Sentences 1.35K 0.27K 0.27K
Entities 1.89K 0.39K 0.42K
OntoNotes Sentences 15.7K 4.3K 4.3K
Entities 13.4K 6.95K 7.7K
CoNLL2003 Sentences 15.0K 3.5K 3.7K
Entities 23.5K 5.9K 5.7K
Table 6: Detail of the two NER datasets.
Appendix B Difficulty Metric Baselines